solve:
cos -1(small in the upper righthand corner)x + cos -1(small in the upper right hand corner) 2x= 60degrees

Answer by stanbon(48558)

solve:
cos -1(small in the upper righthand corner)x + cos -1(small in the upper right hand corner) 2x= 60degrees
----------
cos^-1(x) + cos^-1(2x) = 60 degrees
---------------
Take the cos of both sides to get:
x + 2x = cos(60)
x + 2x = 1/2
3x= 1/2
x = 1/6
--------------
Cheers,
Stan H.